A girl is ten years older than her brother, next year she will be twice as old as him, how old is the girl?
Answer by stanbon(75887)

A girl is ten years older than her brother, next year she will be twice as old as him, how old is the girl?
------
Equations:
g = b + 10
g+1 = 2(b+1)
----
Substitute for "g" and solve for "b":
b+10+1 = 2b+2
b = 9
---
Solve for "g":
g = b + 10 = 19 (girl's age now)
=======
Cheers,
Stan H.
============